OpenStudy (anonymous):

Please Help Solve each literal equation for the stated variable. Equation Posted

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 }iwh,\] solve for i

OpenStudy (anonymous):

i=3V/wh

OpenStudy (anonymous):

since it is all multiplication on the right hand side, just divide by \(\large (\frac{1}{3}\cdot w\cdot h) \) to both sides.

OpenStudy (anonymous):

Is tht the answer The Mind?